A second number is five times a first number. A third number is 100 more than the first number. If the sum of the three numbers is 415, find the number. How do I solve this?
Answer by Earlsdon(6294)

Let the first number be x, then the second number is 5x, and the third number is x+100. The sum of these three numbers is 415.
x+5x+(x+100) = 415 Combine like-terms.
7x+100 = 415 Subtract 100 from both sides.
7x = 315 Divide both sides by 7.
x = 45
5x = 225
x+100 = 145.
Check:
45+225+145 = 415